a bag contains 10 red balls 15 green balls and 5 purple balls if two balls are chosen at random without replacement what is the probability that both of them are the same color?
Answer by stanbon(75887)

a bag contains 10 red balls 15 green balls and 5 purple balls if two balls are chosen at random without replacement what is the probability that both of them are the same color?
-----
# of same-color pairs:: 10C2 + 15C2 + 5C2 = 45+95+10 = 150
------
# of random pairs:: 30C2 = 435
----
P(same-color pills) = 150/435 = 0.345
-------------
Cheers,
Stan H.
-----------